OpenStudy (anonymous):

When you are finding the simplified form of a radical whose radicand contains a number and a variable, what do you do after you simplify the number and the variable separately?

OpenStudy (amistre64):

can you give that in a less abstract arrangment perhpas?

OpenStudy (anonymous):

What?

OpenStudy (amistre64):

for example\[\sqrt{4x^3}=\sqrt{4}\sqrt{x^2}\sqrt{x}=2x\sqrt x\] but im not sure what your question is asking

OpenStudy (anonymous):

I don't have an equation I'm asking how to do something so I know for future refrence

OpenStudy (anonymous):

I guess I can make an equation if you want

OpenStudy (amistre64):

well, when your done simplifying, then your done ... there is nothing left to "do"

OpenStudy (amistre64):

an equation might be useful to practice with, yes

OpenStudy (anonymous):

what do you do after you simplify the number and the variable separately?

OpenStudy (anonymous):

Thats basically what I am asking

OpenStudy (amistre64):

once something is simplified outside of a radical .... it is either worked in with the rest of the setup, or if theres nothing left to do ... then thats it

OpenStudy (amistre64):

\[\frac{\sqrt{4x^3}}{3x}\] \[\frac{2x\sqrt{x}}{3x}\] \[\frac{2\sqrt{x}}{3}\]

OpenStudy (anonymous):

Okay, thanks :3

OpenStudy (amistre64):

youre welcome